# Collective action and entanglement of magnetically active liquid crystal elastomer ribbons

### v1

@@ 2.3 Mathematical model for dynamic entanglement
In additional to topological entanglements, the non-topological adhesion forces themselves play a pivotal role in the assembly of the aggregate (Figure 4d). A representative time-lapse of simulations for ‘non-sticky’ (Supplementary Information) low and high curvature filaments are shown in Figure 4d (see also Suplementary Video section A.1). In this scenario, the filaments behave as smooth objects that are devoid of any cohesive interactions. As a result, their cumulative entanglement is solely due to topological interactions. Importantly, the time evolution of the # of clusters in the ‘non-sticky’ case (Figure 4e) is drastically different than that of the ‘sticky’ case (Figure 4b). At early times (), a very small decrease in the # of clusters is apparent, however both filament types have cluster numbers above 80% of their initial number. The system then maintains an approximately constant rate of neutral collisions which do not lead to entanglement-induced aggregation. This picture is further supported by the excess linking per filament (Figure 4f), which approaches zero at late times, indicating the absence of meaningful aggregation. The fact that the excess linking does not reach zero is due to the choice of initial state (Supplementary Information) and indicates that the process has not fully equilibrated. Notably, the large initial spike in the excess linking per filament for low curvature filaments further demonstrates their higher innate propensity for topological entanglement.
@@
The structure of the clusters formed in simulation can be analyzed in more detail. Following an extended simulation time, low curvature filaments form larger aggregates (Figure 4g) than those formed using high curvature filaments (Figure 4h). The distribution of the number of contacts for low curvature filaments in an aggregate takes larger values than the contact distribution for high curvature filaments. This again illustrates that in large aggregates, low curvature ribbons can form a larger amount of connections, providing a mechanism for aggregate growth. Since the capacity for high curvature filaments to connect saturates faster, continued collisions may result in non-constructive interactions that explain the lower slope in Figure 4c. The above results could be used to further control the self-assembly of clusters with designer geometric and topological properties.
2.4 Mechanical properties of the aggregate
@@
The fluid-like dispersions of ribbons aggregate into solid collectives. The combination of the high aspect ratio of the ribbon geometry and the induced curvature enable the ribbons to entangle and interlock to form physical bonds. The formation of such bonds drives a transition from a liquid dispersion into a solid mass. To confirm the successful formation of a solid phase, oscillatory strain sweep tests of ribbons in silicone oil were carried out using shear rheometry (Figure 5a and Figure 5b). All testing on aggregates was performed following the application of a magnetic field at 72 RPM for 4 minutes from an initially dispersed state ( in Figure 3a). Both storage and loss moduli increase with increased deformation frequency ( aggregates formed at C, Figure 5c). Such a rate dependent response of the material is indicative of viscoelastic response. Control measurements were taken in which no ribbons were present in the oil (black lines) and with the ribbons placed in the original dispersed state without the prior activation of the magnetic field (gray lines), Figure 5d. The oil and the dispersion of unentangled ribbons each behaved as liquids with moduli dramatically lower than the moduli of the entangled aggregates, Figure 5d and Figure 1d. We report the moduli of the dispersion normalized to the entire plate geometry, despite the fact that if an aggregate forms, the region that contains ribbons is smaller than the plates.
@@
The storage modulus of the aggregate depends on the actuated ribbon shape. Namely, ribbons that bend (, blue lines) and ribbons that bend and mildly twist (, green lines) resulted in aggregates with larger stiffness relative to aggregates formed from ribbons with a higher twist per length, in red (cf. Figure 5e and Figure 2). This finding is consistent with measurements taken for LCE ribbon systems formed through static aggregation (i.e., only thermal actuation from an initially overlayed state) performed by Abdelrahman et al. [6]. However, Abdelrahman et al. [6] found purely bending ribbons to form slightly stronger aggregates than those that coil, when formed statically. Our data provide some evidence to support the postulate that coils with a mild twist facilitate better dynamic entanglement by enabling tight assemblies of interlocked ribbons through interweaving rotational motion. A similar mechanism was also identified by Patil et al. [13] for the dynamic entanglement of the California Blackworm.
@@
Heating aggregated samples can increase bond strength within the aggregate. Aggregates of ribbons are formed at C, in the same way as described above. The testing of the aggregate occurs at different temperatures (, Figure 5h, ). We expect that the increase in curvature induced by the thermal actuation above the aggregation temperature will increase the bond strength and possibly also the number of bonds between ribbons. The red line indicates samples that were entangled at C and subsequently cooled down to the crosslinking temperature, C, in which isolated ribbons return to their flat conformation. In large volumes of oil, the aggregate flows and partially redisperses. When the aggregate is placed in the rheometer and allowed to cool, macroscopic flow was not obvious (Supplementary Video section A.1). The cooled samples exhibit a similar storage modulus to that obtained at C with an increased loss modulus throughout the linear viscoelastic region, which implies higher energy dissipation through viscous flow. Samples tested at C have a decreased yield strain as compared to samples tested at C. Increasing temperature after aggregation resulted in changes to aggregate dimensions (Supplementary Figure S4) for which gap heights was appropriately adjusted (section 4 for more details). At C and C, the aggregates possess higher storage moduli and yield stresses. The increases in moduli and yield stress (Figure 5i) signify increases in both aggregate rigidity and bond strength. The data from the C (black line) experiment is only plotted up to the approximate end of the linear viscoelastic region as dissociation was not observed and slipping occurred.
2.5 Disassembly of aggregates
@@
We hypothesized that aggregate disassembly can be induced through magnetic field rotation at high speeds. However, in moderately high viscosity oil ( cP), an entangled aggregate that was subjected to increasingly high rotation speeds of the field did not result in the dissociation of the aggregate, top panel in Figure 6a. Above a critical frequency, magnetic field rotation resulted in vibration of the magnetic domain, indicating that the rotational kinetic energy did not transfer to the ribbon, Supplementary Video section A.1. This loss of synchronicity is due to the critical ‘step-out’ frequency, representing the threshold above which the applied magnetic torque is not strong enough to keep the ribbon synchronized with the field [30]. The step-out frequency can be controlled with the ribbon magnetization, viscosity of the environment, and the external field strength.
@@
The viscosity of the oil surrounding the ribbons is critical to isothermal aggregate dissociation. The average step-out frequency of a single ribbon () at C in a room temperature viscosity of cP oil was found to be approximately 200 RPM, Figure 6b (). However, a 5.6-fold decrease in a room temperature viscosity ( cP) yielded a two-fold increase in the step-out frequency to 400 RPM, Figure 6b (). The step-out frequency was taken as the point in which the response curve departs from the line representing where response frequency equals field rotation frequency (dashed line). When applying increasing field rotation speeds on an entangled aggregate in low viscosity oil ( cP), one minute exposures resulted in an increasing degree of dissociation, bottom panel in Figure 6a. To quantitatively explore the process kinetics, we tracked the evolution of the # of clusters in time over one minute. Similarly to association kinetics, the majority of the transient process ended after approximately 20 s (Supplementary Figure S5), and the # of clusters was averaged over the final 30 second ().
@@
The critical threshold for dissociation depends on the aggregate bond strength. The # of clusters at different magnetic rotation speeds for aggregates formed at C in cP oil and heated to three different temperatures, C is presented in Figure 6c (). The curves indicate that dissociation into clusters becomes substantial only above approximately 200 RPM at C and that the amount of dissociation increases with the increase in supplied rotational energy. An approximate plateau is reached around 400 RPM (i.e., the step-out frequency obtained for cP).
@@
Aggregates entangled at C but tested at increased temperatures show a shift in the magnetic field rotation speed (or activation energy) required for cluster dissociation, as represented by the schematic in Figure 6d. For example, at C, the minimal frequency for dissociation to occur was approximately 320 RPM, and for C there was no effective dissociation, indicating that the critical barrier for dissociation is again higher than the given step-out frequency of 400 RPM.
The relative bond strength between ribbons can be estimated using the activation energy for dissociation. The activation energy, , represents the minimal amount of rotational kinetic energy required to dissociate the physical bonds in the cluster. We can estimate bond strength using a linear regression (Supplementary Figure S6) to an Arrhenius-like model (Supplementary Information A.4). This measure was used to evaluate the critical magnetic field rotation speed, . Aggregates at C, were found to have an activation energy of dissociation nJ (95% CI: nJ), corresponding to RPM (95% CI: RPM). At C aggregates were found to have nJ (95% CI: nJ), corresponding to RPM (95% CI: RPM).
3 Conclusion
We report a method to create macroscopic aggregates capable of autonomous assembly and disassembly. Liquid crystal elastomer ribbons are programmed to transition reversibly from flat to three dimensional shapes combining bending and twisting. A magnetic domain on the ribbon induces motion through an area much larger than the total length of the ribbon in response to an external, rotating, and uniform magnetic field. As the ribbons move, aggregation from dispersed states is observed. Repeated interactions create entangled clusters that grow over time. The aggregates retain their structure when the magnetic field is removed. The governing mechanisms that control these interactions, are studied using a mathematical model that explores their emergent collective topological dynamics. The model supports experimental evidence that the shape of the filaments affects the final size of the aggregate and demonstrates that cohesive forces play a pivotal role in the topological assembly of the aggregate. Namely, lower curvature filaments have a higher propensity for topological entanglement and are capable of forming a larger amounts of interconnections than the tighter, higher curvature filaments. This propensity, however, requires some additional cohesive interaction to be locked-in and facilitate growth. The formed aggregates are shown to exhibit the properties of viscoelastic solids that are able to store and dissipate energy, and the shape of the individual unit is shown to affect both the aggregation dynamics and the properties of the resulting aggregate. Additionally, we induce controlled dissociation through imparting kinetic rotational energy to the individual ribbons at high field rotation speeds. Ribbon shape and the medium in which dissociation occurs govern disassembly. In essence, slower motion of the ribbons enables the constructive entanglement of the ribbons, while higher velocity motion interferes with these processes and results in dissociation. In light of the mechanistic results from the model, the calculated activation energy for dissociation can be assumed to be related to the frictional energy between the aggregated ribbons. The individual motion apparent in both the constructive (assembly) and deconstructive (disassembly) mechanisms are reminiscent of the dynamic behaviors prevalent in aggregating systems in the animal kingdom, such as worms and fire ants. Imparting such abilities into synthetic systems may enable a wide range of materials to be assembled on demand, enabling a range of potential applications from bio-inspired soft robotics to injectable biomaterials.
4 Experimental Section

### v2

@@ RESULTS
Cohesion is required to facilitate aggregate growth through topological interaction. In addition to topological entanglements, the non-topological cohesive forces themselves play a pivotal role in the assembly of the aggregate (Figure 4d). A representative time-lapse of simulations for ‘non-sticky’ (Supplementary Information) low and high curvature filaments are shown in Figure 4d (see also Suplementary Video List of supporting videos). In this scenario, the filaments behave as smooth objects that are devoid of any cohesive interactions. As a result, their cumulative entanglement is solely due to topological interactions. Importantly, the time evolution of the # of clusters in the ‘non-sticky’ case (Figure 4e) is drastically different than that of the ‘sticky’ case (Figure 4b). At early times (), a very small decrease in the # of clusters is apparent, however both filament types have cluster numbers above 80% of their initial number. The system then maintains an approximately constant rate of neutral collisions which do not lead to entanglement-induced aggregation. This picture is further supported by the excess linking per filament (Figure 4f), which approaches zero at late times, indicating the absence of meaningful aggregation. The fact that the excess linking does not reach zero is due to the choice of initial state (Supplementary Information) and indicates that the process has not fully equilibrated. Notably, the large initial spike in the excess linking per filament for low curvature filaments further demonstrates their higher innate propensity for topological entanglement.
@@
Filament curvature controls the resulting aggregate size and shape. Following an extended simulation time, low curvature filaments form larger aggregates (Figure 4g) than those formed using high curvature filaments (Figure 4h). This is further evident from experimental results presented in Figure 3a, and Supplementary Figure S6a. The distribution of the number of contacts for low curvature filaments in an aggregate takes larger values than the contact distribution for high curvature filaments. This again illustrates that in large aggregates, low curvature ribbons can form a larger amount of connections, providing a mechanism for aggregate growth. Qualitative comparison of the resulting simulated aggregates in Figures 4g and Figure 4h and close-up experimental images in Supplementary Figure S6a, shows that the increased curvature of the ribbons results in a tighter entangled core, and less loose or branched edges that limit constructive interactions with new ribbons. Since the capacity for high curvature filaments to connect saturates faster, continued collisions may result in non-constructive interactions that explain the lower slope in Figure 4c. The above results could be used to further control the self-assembly of clusters with designer geometric and topological properties.
Mechanical properties of the aggregate
@@
The fluid-like dispersions of ribbons aggregate into solid collectives. The combination of the high aspect ratio of the ribbon geometry and the induced curvature enable the ribbons to entangle and interlock to form physical bonds. The formation of such bonds drives a transition from a liquid dispersion into a solid mass. To confirm the successful formation of a solid phase, oscillatory strain sweep tests of ribbons in silicone oil were carried out using shear rheometry (Figure 5a and Figure 5b). All testing on aggregates was performed following the application of a magnetic field at 1.2 Hz (72 RPM) for 4 minutes from an initially dispersed state ( in Figure 3a). Both storage and loss moduli increase with increased deformation frequency ( aggregates formed at C, Figure 5c). Such a rate dependent response of the material is indicative of viscoelastic response. Control measurements were taken in which no ribbons were present in the oil (black lines) and with the ribbons placed in the original dispersed state without the prior activation of the magnetic field (gray lines), Figure 5d. The oil and the dispersion of unentangled ribbons each behaved as liquids with moduli dramatically lower than the moduli of the entangled aggregates, Figure 5d and Figure 1d. We report the moduli of the dispersion normalized to the entire plate geometry, despite the fact that if an aggregate forms, the region that contains ribbons is smaller than the plates.
@@
The storage modulus of the aggregate depends on the actuated ribbon shape. Namely, ribbons that bend (, blue lines) and ribbons that bend and mildly twist (, green lines) resulted in aggregates with larger stiffness relative to aggregates formed from ribbons with a higher twist per length, in red (cf. Figure 5e and Figure 2). This finding is consistent with measurements taken for LCE ribbon systems formed through static aggregation (i.e., only thermal actuation from an initially overlayed state) performed by Abdelrahman et al. 6. However, Abdelrahman et al. 6 found purely bending ribbons to form slightly stronger aggregates than those that coil, when formed statically. Our data provide some evidence to support the postulate that coils with a mild twist facilitate better dynamic entanglement by enabling tight assemblies of interlocked ribbons through interweaving rotational motion. A similar mechanism was also identified by Patil et al. 13 for the dynamic entanglement of the California Blackworm.
@@
Heating aggregated samples can increase bond strength within the aggregate. Aggregates of ribbons are formed at C, in the same way as described above. The testing of the aggregate occurs at different temperatures (, Figure 5h, ). The increase in curvature induced by the thermal actuation above the aggregation temperature increases the bond strength and possibly also the number of bonds between ribbons (Supplementary Figure S6). The red line indicates samples that were entangled at C and subsequently cooled down to the crosslinking temperature, C, in which isolated ribbons return to their flat conformation. In large volumes of oil, the aggregate flows and partially redisperses. When the aggregate is placed in the rheometer and allowed to cool, macroscopic flow was not obvious (Supplementary Video List of supporting videos). The cooled samples exhibit a similar storage modulus to that obtained at C with an increased loss modulus throughout the linear viscoelastic region, which implies higher energy dissipation through viscous flow. Samples tested at C have a decreased yield strain as compared to samples tested at C. Increasing temperature after aggregation resulted in changes to aggregate dimensions (Supplementary Figure S8) for which gap heights was appropriately adjusted (section EXPERIMENTAL PROCEDURES for more details). At C and C, the aggregates possess higher storage moduli and yield stresses. The increases in moduli and yield stress (Figure 5i) signify increases in both aggregate rigidity and bond strength. The data from the C (black line) experiment is only plotted up to the approximate end of the linear viscoelastic region as dissociation was not observed and slipping occurred.
Disassembly of aggregates
@@
We hypothesized that aggregate disassembly can be induced through magnetic field rotation at high speeds. However, in moderately high viscosity oil ( Pas), an entangled aggregate that was subjected to increasingly high rotation speeds of the field did not result in the dissociation of the aggregate, top panel in Figure 6a. Above a critical frequency, magnetic field rotation resulted in vibration of the magnetic domain, indicating that the rotational kinetic energy did not transfer to the ribbon, Supplementary Video List of supporting videos. This loss of synchronicity is due to the critical ‘step-out’ frequency, representing the threshold above which the applied magnetic torque is not strong enough to keep the ribbon synchronized with the field 35, 36. The step-out frequency can be controlled through changing the magnetic moment applied on the ribbon, e.g., by changing the ribbon magnetization or the magnitude of the external field strength35, or by changing the viscosity of the environment 36.
@@
The viscosity of the oil surrounding the ribbons is critical to isothermal aggregate dissociation. The average step-out frequency of a single ribbon () at C in a room temperature viscosity of Pas oil was found to be approximately 3.33 Hz (200 RPM), Figure 6b (). However, a 5.6-fold decrease in a room temperature viscosity ( Pas) yielded a two-fold increase in the step-out frequency to 6.67 Hz (400 RPM), Figure 6b (). The step-out frequency was taken as the point in which the response curve departs from the line representing where response frequency equals field rotation frequency (dashed line). When applying increasing field rotation speeds on an entangled aggregate in low viscosity oil ( Pas), one minute exposures resulted in an increasing degree of dissociation, bottom panel in Figure 6a. To quantitatively explore the process kinetics, we tracked the evolution of the # of clusters in time over one minute. Similarly to association kinetics, the majority of the transient process ended after approximately 20 s (Supplementary Figure S9), and the # of clusters was averaged over the final 30 second ().
@@
The critical threshold for dissociation depends on the aggregate bond strength. The # of clusters at different magnetic rotation speeds for aggregates formed at C in Pas oil and heated to three different temperatures, C is presented in Figure 6c (). The curves indicate that dissociation into clusters becomes substantial only above approximately 3.33 Hz (200 RPM) at C and that the amount of dissociation increases with the increase in supplied rotational energy. An approximate plateau is reached around 6.67 Hz (400 RPM) (i.e., the step-out frequency obtained for Pas). Additional estimation of the rotational Reynolds in the new range, taking into account the reduced viscosity and the increased rotation speeds, provides that rotational Reynolds values increase up to approximately, , considered to be close to the approximate limit for laminar flow37, i.e., 40-60.
@@
Aggregates entangled at C but tested at increased temperatures show a shift in the magnetic field rotation speed required for cluster dissociation. This likely represents an increase in the energetic barrier for dissociation, as represented by the schematic in Figure 6d. For example, at C, the minimal frequency for dissociation to occur was approximately 5.3 Hz (320 RPM), and for C there was no effective dissociation, indicating that the critical barrier for dissociation is again higher than provided below the given step-out frequency of 6.67 Hz (400 RPM).
DISCUSSION
We report a method to create macroscopic aggregates capable of autonomous assembly and disassembly. Liquid crystal elastomer ribbons are programmed to transition reversibly from flat to three dimensional shapes combining bending and twisting. A magnetic domain on the ribbon induces motion through an area much larger than the total length of the ribbon in response to an external, rotating, and uniform magnetic field. As the ribbons move, aggregation from dispersed states is observed. Repeated interactions create entangled clusters that grow over time. The aggregates retain their structure when the magnetic field is removed.
The governing mechanisms that control ribbon interactions, are studied using a mathematical model that explores their emergent collective topological dynamics. The model supports experimental evidence that the shape of the filaments affects the final size of the aggregate and demonstrates that cohesive forces play a pivotal role in the topological assembly of the aggregate. Namely, lower curvature filaments have a higher propensity for topological entanglement and are capable of forming larger amounts of interconnections than the tighter, higher curvature filaments. This propensity, however, requires cohesion or friction to avoid the ribbons from slipping past each other as they interact with other ribbons in the suspension. One interpretation may be that the moderate curvature of the ribbons is not sufficient to induce frustration in the ribbons, at least in the early stages of the aggregate formation or at a low number of contacts. The cohesion thus allows the ribbons to be locked-in and facilitate growth.
The resulting aggregates behave as viscoelastic solids. The formed aggregates are shown to exhibit the properties of viscoelastic solids that are able to store and dissipate energy, and the shape of the individual unit is shown to affect both the aggregation dynamics and the properties of the resulting aggregate. Additionally, we induce controlled dissociation through imparting kinetic rotational energy to the individual ribbons at high field rotation speeds. Ribbon shape and the medium in which dissociation occurs govern disassembly. In essence, slower motion of the ribbons enables the constructive entanglement of the ribbons, while higher velocity motion interferes with these processes and results in dissociation. In light of the mechanistic results from the model, the barrier for dissociation can be assumed to be related to the frictional energy between the aggregated ribbons.
The individual motion apparent in both the constructive (assembly) and deconstructive (disassembly) mechanisms are reminiscent of the dynamic behaviors prevalent in aggregating systems in the animal kingdom, such as worms and fire ants. Imparting such abilities into synthetic systems may enable a wide range of materials to be assembled on demand, enabling a range of potential applications from bio-inspired soft robotics to injectable biomaterials.
Limitations of the study and future outlook
This work presents a method for assembling solid aggregates from dilute suspensions of responsive shape-changing polymer ribbons. A previous work by Abdelrahman et al. 6 demonstrated the assembly of similar polymer ribbons from concentrated suspensions by changing temperature. This temperature change induces morphing in the individual ribbons from flat to bent and twisted shapes. These curvatures form topological interactions that act as physical bonds that provide structural stability to the aggregate, providing it with solid-like properties. However, since motion arises solely due to the actuation of the ribbon, such interactions require close proximity between the ribbons (smaller than half a ribbon length), i.e., high concentrations. Herein, we added an additional mechanism for motion, induced by an external, rotating magnetic field. This additional motion allows the separation of the mechanism that controls the shape of a ribbon from the mechanism that drives the interaction of the individual ribbons. These separate mechanisms enable aggregates to form with solid-like properties at constant temperatures and low starting concentrations. Fundamentally, this achievement overcomes the limitation on distance between ribbons to lengths much larger than the length of a single unit 6. However, we note that the current method is limited to the region in which the field is uniform (approximately 60 mm diameter circle, Supplementary Figure S3). Such limitations may prevail through other external sources of propulsion, whereas internal propelling forces 38, 39, 40 may completely eliminate this limitation. Future studies may seek to study the effects of such alternative mechanisms.
Magnetic-field induced motion opens the way for advanced system manipulation. Imparting the individual ribbons with magnetic responsiveness allows for variation in ribbon response by modulation of the magnetic field. This work has demonstrated that change in the rotation frequency allows for transition between assembly and disassembly dynamics between ribbons. Additional progressions of this work may include the individual control of different ribbon populations that have different magnetization values, resulting in different step-out frequencies and different dipole moments or magnetic torques under an applied field. For example, Nelson et al. 35 demonstrated that systems of rotating magnetic microrobots can be controlled simultaneously when operated above the step-out frequency. Wie et al. 41 used modulation of the magnetic field that induce centrifugal forces to control micro-robots, an approach that has the potential to contribute additional control to the ribbon dynamics. Such concepts, along with careful design of the geometry of the magnetically responsive unit and modal manipulations of the applied magnetic field, have been shown to result in variable collective task-execution and self-organization 42, 36, 26.
The assembly of aggregates in a fluid medium requires a better understanding of hydrodynamic effects. This work has demonstrated the important contribution of the fluid medium on the dynamics of individual ribbon motion and disassembly of the collective. Expectantly, the induced motion of the individual ribbons and ribbon clusters induces flow. This was briefly demonstrated in this work using simple flow visualization in different scenarios (Supplementary Video List of supporting videos). This visualization allows us to eliminate the occurrence of large-scale circulatory flows that act as sinks, resulting in ribbon clustering. Additionally, the fluid does not move the ribbons once the field is stopped. Therefore, we have postulated that under the assembly conditions of this study, hydrodynamics play a lesser role as compared to ribbon form. There has been significant discussion in the literature on the effect of a compliant filament in different flow fields 43, 44. Additionally, high rotation speeds may result in local turbulence that can affect the dynamics of dissociation. Future studies may choose to employ flow measurement techniques to elucidate such phenomena.
Linking between ribbons relies on cohesive interactions. The mathematical model presented in this work demonstrated that the successful formation of topological links between curved ribbons relies on the existence of forces between the surfaces of the ribbons. The contact force law considered here is a model for adhesion, where contacts become pinned until they are separated by a sufficiently large force. However, there are several other examples of contact laws that incorporate other effects, such as sliding contacts, which are not captured here. Understanding how aggregation is affected by these different contact force protocols, such as static friction, represents an important future challenge. It may also be inferred that for aggregate dissociation to occur, such cohesive interactions must be overcome. However, this study did not discern the energy of these interactions and the effect its modulation would exert on the resulting aggregate shape and size. Quantitative information on the nature of interactions within the aggregates requires three-dimensional imaging capabilities such as micro-CT 6 or ultrasound 13. Future work may explore a more in-depth characterization and estimation of the work required to separate interlocked ribbons, for example as a function of variable surface roughness or contact surface area.
Ribbon compliance and dimensions can significantly affect the behavior of the system. The use of polymer ribbons for aggregation as presented in this study exists in a vast parameter space. Two important parameters that have been fixed in this study were the crosslink density (affecting its elastic modulus) and the dimensions of the ribbons. These parameters act to change the effective ribbon stiffness. Namely, in this work we studied the ribbons in the rigid limit, where the magnetic field induces mostly rigid body rotations rather the bending moments. Relaxing this limitation, with lower modulus or longer/thinner ribbons, is expected to substantially change the mechanisms of interaction, the importance of hydrodynamic effects, and the resulting aggregate properties, and mechanical response. Future studies realizing such flexible and highly compliant filaments are expected to unlock unique behaviors.
EXPERIMENTAL PROCEDURES